The current price of AA stock is $60. Dividends are expected to grow at 5% indefinitely and the most recent dividend (D0) was $3. What is the required rate of return on AA stock?
Answer #1

Price = D1/(ke-g) = $60.00 D1 = $3.15 3*1.05 5.00% ke - g = 5.25% (3.15/60 ] required return = ke 10.25% [5% +5.25%]